The balance sheet for Levy Corp. is shown here in market value terms. There are 8,000 shares of stock outstanding.

  

Market Value Balance Sheet
  Cash $ 44,000   Equity $ 384,000
  Fixed assets 340,000
     Total $ 384,000      Total $ 384,000

  

The company has declared a dividend of $1.70 per share. The stock goes ex-dividend tomorrow.

Ignoring any tax effects, what is the stock selling for today? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))

  

  Current stock price $  per share  

  

Ignoring any tax effects, what will it sell for tomorrow? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))

  

  New stock price $  per share  

  

Ignoring any tax effects, what will the balance sheet look like after the dividends are paid? (Do not round intermediate calculations.)
Market Value Balance Sheet
  Cash $      Equity $   
  Fixed assets   
     Total $         Total $   
Answer #1

1- Current Value of share= Total market value of equity

No. of Share

=$384000/8000=$48 per share

2.- Total Dividend=8000*1.70$=$13600

Cash Balance=$44000-13600=$30400

Market Value Balance Sheet
  Cash $ 30,400   Equity $ 370,400
  Fixed assets 340,000
     Total $ 370,400      Total $ 370,400

Price per Share=370400/8000=$46.3
Answer #2

(1)

Current stock price = 384,000/8,000 = 48

(2)

New stock price = 48 - 1.7 = 46.3

(3)

Market Value Balance Sheet
  Cash $

30,400

=44,000-8,000*1.7

  Equity $ 370,400
  Fixed assets

340,000

     Total $

370,400

     Total $ 370,400